OpenStudy (meilendurcer):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A. Y= 4over3x + 8 B. Y= 4over3x + 3 C. Y= 3over4x + 8 D. Y= 3over4x + 3 Please help again.

Miracrown (miracrown):

What do you think?

OpenStudy (anonymous):

if the lines are parallel then the slopes will be same....so get the slope from the 1st equation....and equation of line will be like y-y1 = m(x-x1) where x1, y1 are the points through which the line is passing

OpenStudy (meilendurcer):

sorry that took a long time, i was on mobile but i think the answer is C.

OpenStudy (anonymous):

cant be C...slopes are not equal

OpenStudy (meilendurcer):

true

OpenStudy (meilendurcer):

my first answer was B but maybe A, i'm really not sure

OpenStudy (anonymous):

its A

OpenStudy (meilendurcer):

Great thanks! can you help with another?

OpenStudy (anonymous):

yeah sure..

OpenStudy (meilendurcer):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = 3over2sx - 1 y = 3over2x + 3 y = -2over3x - 1 y = -2over3x + 3

OpenStudy (meilendurcer):

I know it's not A

OpenStudy (meilendurcer):

I'm leaning near the positive side but it has -6 so i can't tell.

OpenStudy (meilendurcer):

is it C?

OpenStudy (anonymous):

its C

OpenStudy (meilendurcer):

if you can, can i have help with some others?

OpenStudy (anonymous):

yep :)

OpenStudy (meilendurcer):

k cool. This one has a graph so i'll attatch it. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3, 2). y = 1 over 3 x + 3 y = -1 over 3x + 3 y = 3 x + 2 y = - 1 over 3 x + 2

OpenStudy (meilendurcer):

would it be C or D? i'm thinking those.

OpenStudy (anonymous):

its B

OpenStudy (meilendurcer):

how did you get that?

OpenStudy (anonymous):

coz y-2 =-1/3x +1

OpenStudy (meilendurcer):

ahh ok i see.

OpenStudy (meilendurcer):

Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 My answer is B y = 3. I know it's not A.

OpenStudy (anonymous):

yep its b

OpenStudy (meilendurcer):

really? great!

OpenStudy (anonymous):

yep ;)